Selected Answer o [None Given] Comed Answer 130 N uestion 14 Newtons on the Moon? (HINT Moon g 163 m/s 2)
Answer #1

gravitational force on the moon=m*g1

here g1=acceleration due to gravity on the surface of Moon.

weight on the moon= 80*1.63=130.4N